OpenStudy (anonymous):

If g(x) = x2 + 2, find g(3). 9 8 11 6

OpenStudy (anonymous):

11

OpenStudy (anonymous):

All you have to do is plug in 3 for x.

OpenStudy (anonymous):

then how is it 11?

OpenStudy (anonymous):

g(3)=(3)2+2 g(3)=6+2 g(3)=8 the answer is not 11

OpenStudy (anonymous):

is it 4? @Rosa34

OpenStudy (anonymous):

2

OpenStudy (jhannybean):

First off,stop guessing answers, it's not going to help :\

OpenStudy (anonymous):

No @shahbaaz786 it is 8

OpenStudy (anonymous):

you are done solving after that step

OpenStudy (anonymous):

isnt ir if g=2 THEN 2X2X2 EQUALS 8 RIGHT ?

OpenStudy (jhannybean):

What @Rosa34 is saying is \(\large g(x) = x^2 +2\) and \(\large x=3\) so wherever you see an x in your equation for \(g(x)\) you plug in 3.

OpenStudy (jhannybean):

\[\large g(3) = (3)^2 +2 = 9 +2 = 11\]

OpenStudy (jhannybean):

x2 does not mean 2x.

OpenStudy (anonymous):

@Jhannybean are you giving him an example?

OpenStudy (jhannybean):

Nope. He lazily copy pasted the problem. Its supposed to say \(x^2\) but when you copy paste over the question it comes out as x2. If it really was 2(x)he would have written it 2x and not x2.

OpenStudy (anonymous):

@Jhannybean that makes since i just thought he wrote it backwards. Nice looking out.

OpenStudy (jhannybean):

No problem

OpenStudy (anonymous):

so then the answer wrong

OpenStudy (anonymous):

?

OpenStudy (jhannybean):

Your answer is not 4, or 8.

OpenStudy (anonymous):

so its 6 ?

OpenStudy (anonymous):

@shahbaaz786 yes that answer is wrong because it was the wrong equation. is that the equation you want me to answer If g(x) = x2 + 2, find g(3).

OpenStudy (anonymous):

so is it 6?

OpenStudy (jhannybean):

We're not here to give you answers. You can follow what we've done in the posts above to figure out how to solve the problem.

OpenStudy (anonymous):

Ok i'm confused now what equation are we solving?

OpenStudy (anonymous):

i did on my paper and i got 6

OpenStudy (jhannybean):

Can you show us how you got 6?

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

@Jhannybean is right can u type how you solved it and we'll tell you if you did the work right or not.

OpenStudy (anonymous):

how i don't know the right keys on the mac

OpenStudy (anonymous):

r u guys gonna reply ?

OpenStudy (jhannybean):

just type up what you wrote in the text box :)

OpenStudy (anonymous):

its too complicated ill just guess

OpenStudy (anonymous):

@shahbaaz786 your being complicted sk someone to help you at home.

OpenStudy (anonymous):

ask

OpenStudy (anonymous):

their working

OpenStudy (anonymous):

so its the same equation @Jhannybean typed above?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok so i cant read that.

OpenStudy (anonymous):

click on it ?

OpenStudy (anonymous):

ok ya the answer is correct, but do you now how you got it.

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok

OpenStudy (jhannybean):

The answer, once again, is not 8. \[\large (3)^2 +2 = 9 + 2 = 11\]